# On Reducing the Execution Latency of Superconducting Quantum Processors via Quantum Program Scheduling

### v1

@@ 4.5. Sensitivity Analysis
We study the impact of four hyperparameters, i.e. width weight $\alpha$, shot weight $\beta$, time weight $\gamma$, and maximum usage $\eta$. The sensitivity tests for $\alpha$, $\beta$, and $\gamma$ are implemented on the noise model, while the test for $\eta$ is done on Xiaohong, because $\eta$ is more related with PST, and PST on Xiaohong reflects the realistic effect on QPUs.
Time weight $\gamma$. As shown in Fig. 4b, with the increase of $\gamma$, the standard deviation of TAT is getting smaller while average TAT is getting longer. Hence, there is a trade-off between fairness and time metrics. Since time metrics are more important for both users and suppliers, we pick $\gamma=0.5$ in our experiments.
Width weight $\alpha$ and shot weight $\beta$. $\alpha$ and $\beta$ are more concerned with TAT and QPU time. As shown in Fig. 4c and Fig. 4d, the increase in $\alpha$ leads to decline in average TAT and rise in QPU time. By contrast, larger $\beta$ reduces QPU time, and average TAT reaches its minimum when $\beta=2$. To strike a balance between average TAT and QPU time, we pick $\alpha=3,\beta=2$ for the noise model.
@@
Maximum usage $\eta$. It directly influences TRF. As shown in Fig. 4a, with its increase, both average TAT and PST decline. Because more programs are executed in parallel. Hence, we can balance the fidelity and execution latency by tuning $\eta$ and we set $\eta=5/6$.
In general, our method is robust to these hyperparameters.
4.6. Ablation Study of $*{}^{*}$
@@
To evaluate our $EPST^{*}$ score, we execute the candidate circuits through the serial execution mode with and without $EPST^{*}$. For the noise model, $EPST^{*}$ raises PST by 1.59% (from 72.22% to 73.81%). For Xiaohong, it raises PST by 0.94% (from 17.73% to 18.67%). Low PST on Xiaohong is due to its fluctuating noise. Hence, $EPST^{*}$ can improve the fidelity of quantum programs.
5. Conclusion
@@
We have formulated the Quantum Program Scheduling Problem to boost the execution efficiency of (superconducting) quantum processors. Our scheduling method perceives the impact of different programs on time metrics through our priority score, and the noise-aware initial mapping improves the fidelity. Also, three greedy baselines are given for comparison. Results show that our method outperforms greedy baselines in both QPU time and TAT. Besides, the fidelity and fairness are also guaranteed. The small runtime overhead shows its scalability on large QPUs. We envision that our scheduling method may be further adapted to non-superconducting quantum cloud, which we leave for future work.
Acknowledgements.
This work is partly supported by the QuantumCTek Quantum Computing Cloud Platform and NSFC (92370201).

### v2

@@ 4.7. Sensitivity Analysis of Hyperparameters
We study the impact of five hyperparameters: width weight $\alpha$, shot weight $\beta$, time weight $\gamma$, maximum usage $\eta$, and time interval $\Delta t$. The sensitivity tests for $\alpha$, $\beta$, $\gamma$, and $\Delta t$ are implemented on the noise model; the test for $\eta$ is done on Xiaohong, as $\eta$ is more related with PST, and PST on Xiaohong reflects the realistic effect on QPUs.
@@
Maximum usage $\eta$. $\eta$ directly influences TRF and the qubit utilization rate. With its increase, both average TAT and PST decline as shown in Fig. 6a, because more jobs are executed in parallel. Hence, we can balance the fidelity and execution latency by tuning $\eta$ and we set $\eta=5/6$.
Time weight $\gamma$. As shown in Fig. 6b, with the increase of $\gamma$, the standard deviation of TAT is getting smaller while average TAT is getting longer. Hence, there is a trade-off between fairness and time metrics. Since time metrics are more important for both users and suppliers, we pick $\gamma=1.0$ in our experiments.
Width weight $\alpha$ and shot weight $\beta$. $\alpha$ and $\beta$ are more concerned with TAT and QPU time. As shown in Fig. 7a and Fig. 7b, the increase in $\alpha$ leads to decline in average TAT and rise in QPU time. By contrast, larger $\beta$ reduces QPU time, and average TAT reaches its minimum when $\beta=4.5$. To strike a balance between average TAT and QPU time, we pick $\alpha=6,\beta=4.5$ for the noise model.
Time interval $\Delta t$. We introduce $\Delta t$ to avoid the starvation problem. $\Delta t$ should be big enough, or it will result in FIFO-p manner execution. As shown in Fig. 6c, with the increase of $\Delta t$, average TAT decreases while maximum TAT rises, indicating that quantum jobs with large width or shot numbers waits longer to be executed. Nevertheless, the average TAT and maximum TAT saturates at $\Delta t=360$. Therefore, we set $\Delta t=360$ to avoid the starvation problem and guarantee low average TAT.
4.8. Ablation Study of EPST*
@@
To evaluate our $EPST^{*}$ score, we run NAQJS with and without $EPST^{*}$. For the noise model, $EPST^{*}$ raises PST by 0.70% (from 68.95% to 69.65%). For Xiaohong, it raises PST by 1.77% (from 33.93% to 35.70%). Hence, $EPST^{*}$ can improve the fidelity of quantum circuits.
5. Conclusion and Future Work
@@
We have formulated the Quantum Job Scheduling Problem (QJSP) and proposed a noise-aware quantum job scheduler (NAQJS) to boost the execution efficiency of (superconducting) quantum processors. Our scheduling method perceives the impact of different jobs on time metrics through our priority score, and the noise-aware initial mapping improves the fidelity. Results show that NAQJS outperforms all the baselines in both QPU time and TAT. Besides, the fidelity and fairness are also guaranteed. The small runtime overhead shows its scalability on large QPUs. QJSP may be more important when noise-free quantum processors emerges, because it will directly influence the efficiency of them.
Future Work. In this paper, we conduct experiments on a 66-qubit quantum processor, which is the largest in scale compared with related works. Since larger-scale QPUs are not yet open to the public or too expensive to use, we leave larger-scale experiments for our future work. Besides, we envision that NAQJS may be further adapted to non-superconducting quantum cloud.